• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name hsmith

    Log entry time 15:21:10 on July 06, 2010

    Entry number 330479

    keyword=Left Q1

    The left q1 is up and the current is stable at 1.131 Gev. To take cabling issues out of the equation the control card extension cabling has been removed. However, this did not alleviate the oscillations. We replaced the DAC with the spare and the magnet miss-match issue went away. Additionally, the output set point and the read back were better aligned. The current fluctuations are now within acceptable tolerances. Will continue to monitor.